# CreatiLayout: Siamese Multimodal Diffusion Transformer for Creative Layout-to-Image Generation

### v1

@@ Abstract
Corresponding author.
Diffusion models have been recognized for their ability to generate images that are not only visually appealing but also of high artistic quality. As a result, Layout-to-Image (L2I) generation has been proposed to leverage region-specific positions and descriptions to enable more precise and controllable generation. However, previous methods primarily focus on UNet-based models (e.g., SD1.5 and SDXL), and limited effort has explored Multimodal Diffusion Transformers (MM-DiTs), which have demonstrated powerful image generation capabilities. Enabling MM-DiT for layout-to-image generation seems straightforward but is challenging due to the complexity of how layout is introduced, integrated, and balanced among multiple modalities. To this end, we explore various network variants to efficiently incorporate layout guidance into MM-DiT, and ultimately present SiamLayout. To Inherit the advantages of MM-DiT, we use a separate set of network weights to process the layout, treating it as equally important as the image and text modalities. Meanwhile, to alleviate the competition among modalities, we decouple the image-layout interaction into a siamese branch alongside the image-text one and fuse them in the later stage. Moreover, we contribute a large-scale layout dataset, named LayoutSAM, which includes 2.7 million image-text pairs and 10.7 million entities. Each entity is annotated with a bounding box and a detailed description. We further construct the LayoutSAM-Eval benchmark as a comprehensive tool for evaluating the L2I generation quality. Finally, we introduce the Layout Designer, which taps into the potential of large language models in layout planning, transforming them into experts in layout generation and optimization. Our code, model, and dataset will be available at Project Page.
Figure 1: We present a novel approach to empower MM-DiT for layout-to-image generation, achieving high-quality and fine-grained controllable generation, as evidenced by the precise rendering of complex attributes (e.g. color, texture, shape, numeracy, and text).
@@ 2 Related Work
To achieve more precise and controllable generation, layout-to-image generation [37, 62, 72, 12, 8, 14, 57, 47, 34, 65, 27, 68, 71, 67, 21] has been proposed to generate images based on layout guidance, which includes several entities. Each entity comprises a spatial location and a region description. However, previous methods primarily focused on UNet-based architectures, and enabling MM-DiT for layout-to-image generation is challenging due to the complexity of introducing, integrating, and balancing the layout among multiple modalities. In this paper, we focus on exploring winner solutions for incorporating layout into MM-DiT, thereby fully unleashing its capabilities and achieving precisely controllable generation.
@@ 3.2 Layout-to-Image Generation
where [·, ·] denotes concatenation across the tokens dimension, $\mathbf{Q}^{z}=\bm{h}^{z}\mathbf{W}_{q}^{z}$, $\mathbf{K}^{p}=\bm{h}^{z}\mathbf{W}_{k}^{z}$, $\mathbf{V}^{z}=\bm{h}^{z}\mathbf{W}_{v}^{z}$; and $\mathbf{W}_{q}^{z}$, $\mathbf{W}_{k}^{z}$, $\mathbf{W}_{v}^{z}$ are the weight matrices for the query, key, and value linear projection layers for image tokens, respectively. Tokens $\bm{h}^{p}$ of the global caption $\bm{p}$ are handled by the same paradigm as $\bm{h}^{z}$ but with their own weights. ${\bm{h}^{z}}^{\prime}$ and ${\bm{h}^{p}}^{\prime}$ are the image and caption tokens after interaction. To this end, the next critical step is to incorporate the layout tokens $\bm{h}^{l}$. We explore three variants of network designs that incorporate layout tokens, as shown in Fig. 2.
@@
In this paper, SiamLayout is chosen as the network that incorporates layout into MM-DiT, and our primary experiments are conducted on it.
@@
where ${\lambda}_{region}$ modulates the importance of $\mathcal{L}_{region}$. During the inference phase, we perform layout-conditioned denoising only in the first 30% of the steps [37].
@@ 4.1 Experimental Details
We conduct experiments on two types of datasets: the fine-grained open-set LayoutSAM and the coarse-grained closed-set COCO 2017 [40]. For LayoutSAM, we train on 2.7 million image-text pairs with 10.7 million entities and conduct evaluations on LayoutSAM-Eval, which includes 5,000 prompts with detailed entity annotations. We measure generation quality using the metrics outlined in Sec. 3.3. For COCO, following previous work [71, 68], we filter out bounding boxes smaller than 2% of the total image area and images containing dense bounding boxes and crowds, resulting in 61,002 training images and 2,565 validation images. We use the YOLO-v11-x [28] to validate the model’s layout adherence by detecting objects in the generated images and then calculating AP, $\mathrm{AP}^{50}$, and AR against the ground truth. We use FID, CLIP score, and IS to measure the global image quality.
@@
We employ experiments on SD3-medium, and the extra parameters introduced by SiamLayout amount to 1.28B. The training resolution for the LayoutSAM dataset is $1024\times 1024$, and $512\times 512$ for COCO. We utilize the AdamW optimizer with a fixed learning rate of 5e-5 and train the model for 600,000 iterations with a batch size of 16. We train SiamLayout with 8 A800-40G GPUs for 7 days. The value of $\lambda_{region}$ is set to 2. LayoutDesigner is fine-tuned on Llama-3.1-8B-Instruct [9] for one day using one A800-40G GPU.
@@
Spatial $\uparrow$ Color $\uparrow$ Texture $\uparrow$ Shape $\uparrow$ IR $\uparrow$ [66] Pick $\uparrow$ [31] CLIP $\uparrow$ FID $\downarrow$ IS $\uparrow$
@@
SiamLayout 92.67 74.45 77.21 75.93 69.47 22.02 34.01 19.10 22.04
vs. prev. SoTA +4.68 +5.29 +4.43 +4.85 +60.33 +1.01 +1.59 +0.57 +1.47
Table 2: Quantitative comparison on the LayoutSAM-Eval. Bold and underline represent the best and second best methods, respectively.
@@ 4.2 Evaluation on Layout-to-Image Generation
Figure 5: Qualitative results on LayoutSAM-Eval. SiamLayout outperforms previous methods by a clear margin in generating entities with complex attributes like color, texture, text, and portraits.
@@ 4.5 Ablation Study
Figure 7: Qualitative comparison of different layout planners.
@@ 5 Conclusion
A More details on Datasets and Benchmarks
A.1 LayoutSAM Dataset and Benchmark
@@
A.2 Layout Planning Dataset and Benchmark
@@
B More analysis on modal competition
@@
C More qualitative results
We present more qualitative results in Fig. 13. Experimental results show that our proposed method empowers MM-DiT for layout-to-image generation, achieving visually appealing and precisely controllable generation, as demonstrated by the high adherence to complex attributes such as color, texture, and shape.

### v2

@@ Abstract
Diffusion models have been recognized for their ability to generate images that are not only visually appealing but also of high artistic quality. As a result, Layout-to-Image (L2I) generation has been proposed to leverage region-specific positions and descriptions to enable more precise and controllable generation. However, previous methods primarily focus on UNet-based models (e.g., SD1.5 and SDXL), and limited effort has explored Multimodal Diffusion Transformers (MM-DiTs), which have demonstrated powerful image generation capabilities. Enabling MM-DiT for layout-to-image generation seems straightforward but is challenging due to the complexity of how layout is introduced, integrated, and balanced among multiple modalities. To this end, we explore various network variants to efficiently incorporate layout guidance into MM-DiT, and ultimately present SiamLayout. To Inherit the advantages of MM-DiT, we use a separate set of network weights to process the layout, treating it as equally important as the image and text modalities. Meanwhile, to alleviate the competition among modalities, we decouple the image-layout interaction into a siamese branch alongside the image-text one and fuse them in the later stage. Moreover, we contribute a large-scale layout dataset, named LayoutSAM, which includes 2.7 million image-text pairs and 10.7 million entities. Each entity is annotated with a bounding box and a detailed description. We further construct the LayoutSAM-Eval benchmark as a comprehensive tool for evaluating the L2I generation quality. Finally, we introduce the Layout Designer, which taps into the potential of large language models in layout planning, transforming them into experts in layout generation and optimization.
Figure 1: We present a novel approach to empower MM-DiTs for layout-to-image generation, achieving high-quality and fine-grained controllable generation, as evidenced by the precise rendering of complex attributes (e.g. color, texture, shape, and text).
@@ 2 Related Work
To achieve more precise and controllable generation, layout-to-image generation [38, 64, 74, 13, 9, 15, 59, 49, 35, 67, 28, 70, 73, 69, 22, 44, 7] has been proposed to generate images based on layout guidance, which includes several entities. Each entity comprises a spatial location and a region description. However, previous methods primarily focused on UNet-based architectures, and enabling MM-DiT for layout-to-image generation is challenging due to the complexity of introducing, integrating, and balancing the layout among multiple modalities. In this paper, we focus on exploring winner solutions for incorporating layout into MM-DiT, thereby fully unleashing its capabilities and achieving precisely controllable generation.
@@ 3.2 Layout-to-Image Generation
where [·, ·] denotes concatenation across the tokens dimension, $\mathbf{Q}^{z}=\bm{h}^{z}\mathbf{W}_{q}^{z}$, $\mathbf{K}^{p}=\bm{h}^{z}\mathbf{W}_{k}^{z}$, $\mathbf{V}^{z}=\bm{h}^{z}\mathbf{W}_{v}^{z}$; and $\mathbf{W}_{q}^{z}$, $\mathbf{W}_{k}^{z}$, $\mathbf{W}_{v}^{z}$ are the weight matrices for the query, key, and value linear projection layers for image tokens, respectively. Tokens $\bm{h}^{p}$ of the global caption $\bm{p}$ are handled by the same paradigm as $\bm{h}^{z}$ but with their own weights. ${\bm{h}^{z}}^{\prime}$ and ${\bm{h}^{p}}^{\prime}$ are the image and caption tokens after interaction. To this end, the next critical step is to incorporate the layout tokens $\bm{h}^{l}$. We explore three variants of network designs that incorporate layout tokens, as shown in Fig. 2. Here, we present an illustration of our method based on the SD3 model. For further details regarding FLUX’s layout control mechanism, please refer to the supplementary material.
@@
In this paper, SiamLayout is chosen as the network that incorporates layout into MM-DiTs, and our primary experiments are conducted on it. In addition, we develop a LoRA-based variant of SiamLayout, which achieves comparable layout control accuracy in a more lightweight manner. Further details can be found in the supplementary material.
@@
where ${\lambda}_{region}$ modulates the importance of $\mathcal{L}_{region}$. During the inference phase, we perform layout-conditioned denoising only in the first 30% of the steps [38, 74].
@@ 4.1 Experimental Details
We conduct experiments on two types of datasets: the fine-grained open-set LayoutSAM and the coarse-grained closed-set COCO 2017 [41]. For LayoutSAM, we train on 2.7 million image-text pairs with 10.7 million entities and conduct evaluations on LayoutSAM-Eval, which includes 5,000 prompts with detailed entity annotations. We measure generation quality using the metrics outlined in Sec. 3.3. For COCO, following previous work [73, 70], we filter out bounding boxes smaller than 2% of the total image area and images containing dense bounding boxes and crowds, resulting in 61,002 training images and 2,565 validation images. We use the YOLO-v11-x [29] to validate the model’s layout adherence by detecting objects in the generated images and then calculating AP, $\mathrm{AP}^{50}$, and AR against the ground truth. We use IR [68], Pick [32], FID, CLIP, and IS to measure the global quality.
@@
We employ experiments on SD3-medium and FLUX.1-dev. The training resolution for the LayoutSAM dataset is $1024\times 1024$, and $512\times 512$ for COCO. We utilize the AdamW optimizer with a fixed learning rate of 5e-5 and train the model for 600,000 iterations with a batch size of 16. We train SiamLayout with 8 A800-40G GPUs for 7 days. The value of $\lambda_{region}$ is set to 2. LayoutDesigner is fine-tuned on Llama-3.1-8B-Instruct [10] for one day using one A800-40G GPU.
@@
Spatial $\uparrow$ Color $\uparrow$ Texture $\uparrow$ Shape $\uparrow$ IR $\uparrow$ Pick $\uparrow$ CLIP $\uparrow$ FID $\downarrow$ IS $\uparrow$
@@
HiCo [44] 87.04 69.19 72.36 71.10 12.36 21.70 32.18 22.61 20.15
SiamLayout-SD3 92.67 74.45 77.21 75.93 69.47 22.02 34.01 19.10 22.04
SiamLayout-FLUX 95.67 80.71 83.53 82.80 80.48 22.16 33.92 16.12 21.81
vs. prev. SoTA +7.68 +11.52 +10.75 +11.70 +68.12 +0.46 +1.59 +0.57 +1.47
Table 2: Quantitative results on the LayoutSAM-Eval. Bold, underline, and double underline represent the best, second, and third best methods, respectively.
@@ 4.2 Evaluation on Layout-to-Image Generation
Figure 5: Qualitative results on the LayoutSAM-Eval. SiamLayout outperforms previous methods by a clear margin, especially in generating complex attributes such as color, texture, text, and portrait details. Zoom in for details.
@@ 4.3 Evaluation on Text-to-Image Generation
4.4 Evaluation on Layout Planning
Figure 6: Qualitative comparison of different layout planners.
@@ 5 Conclusion
A More Details on SiamLayout-FLUX
Figure 9: An overview of SiamLayout-FLUX.
FLUX is another outstanding text-to-image generative model based on the Multimodal Diffusion Transformer (MM-DiT) architecture, demonstrating remarkable performance alongside SD3. The core components of FLUX include the double-stream block and the single-stream block. To further verify the generality of our proposed SiamLayout approach, we integrate SiamLayout into FLUX and propose SiamLayout-FLUX, aiming to empower FLUX for layout-to-image generation.
Specifically, as illustrated in Fig. 9, for the Double Stream Block in FLUX, we adopt the same layout integration strategy as presented in Fig. 2. For the Single Stream Block, we maintain the core concept from SiamLayout, which emphasizes preserving the advantage of MM-Attention in facilitating effective multimodal interactions while reducing interference and competition among modalities. Therefore, we similarly introduce a Siamese branch dedicated to processing the interactions between images and layouts. Such a design enables the layout information to independently and concurrently guide the image content generation, analogous to the role played by the global text caption.
Both quantitative and qualitative experimental results demonstrate that our proposed SiamLayout-FLUX outperforms previous methods by a clear margin, validating the effectiveness of SiamLayout and its general applicability across different MM-DiTs.
B LoRA Version of SiamLayout
Figure 10: An overview of SiamLayout-LoRA.
We further propose a more lightweight variant for layout control, named SiamLayout-LoRA, which achieves comparable layout control accuracy with fewer additional parameters. As illustrated in Fig. 10, SiamLayout-LoRA integrates Low-Rank Adaptation (LoRA) modules into the Linear1, Linear2, and MLP1 layers, efficiently handling image-layout interactions based on the frozen pre-trained weights. This LoRA-based variant provides alternative trade-offs between extra parameter overhead and layout control accuracy. In our experiments, we set the LoRA rank to 256.
C Discussion on Extra Parameters and Computation Introduced by Layout Control
Extra Params $\downarrow$ Extra MACs $\downarrow$ Spatial $\uparrow$ Color $\uparrow$ Texture $\uparrow$ Shape $\uparrow$
GLIGEN 24.20% 30.80% 77.53 49.41 55.29 52.72
InstanceDiff 42.67% 155.83% 87.99 69.16 72.78 71.08
HiCo 42.03% 263.70% 87.04 69.19 72.36 71.1
Layout Adapter 16.90% 21.70% 88.43 71.67 73.56 72.61
$\mathrm{M}^{3}$-Attention 48.50% 15.70% 79.15 60.19 62.96 61.29
SiamLayout-SD3 62.40% 38.70% 92.67 74.45 77.21 75.93
SiamLayout-SD3-LoRA 15.90% 39.90% 89.71 71.89 74.14 72.63
Table 7: Comparison of extra parameters and computation. Bold, underline represent the best and second best methods, respectively.
As shown in Tab. 7, we compare different approaches with respect to extra parameters and computational overhead (i.e., MACs required for one denoising step given a layout containing 10 entities). To ensure a fair comparison, we report the relative increase in parameters and computation introduced by integrating layout control compared to the base model.
Our proposed approach outperforms others at comparable additional parameter counts while requiring significantly lower extra computational overhead compared to InstanceDiff and HiCo. This is due to InstanceDiff and HiCo computing image-layout attention separately for each entity before integrating them, causing their computational overhead to increase linearly as the number of entities grows. In addition, our proposed LoRA variant achieves comparable layout control accuracy with substantially fewer extra parameters.
D User Study
Figure 11: The results of the user study.
To appropriately design the evaluation criteria for our user study, we solicited feedback from professional designers as well as a large number of general users. As shown in Fig. 11, the evaluation dimensions were eventually categorized into two groups: region-wise quality and global-wise quality. Specifically, region-wise quality assesses the accuracy of attributes such as spatial, color, shape, texture, text, and portrait details, while global quality evaluates the overall aesthetic quality, prompt-following capability, and reasonability of the generated images. We randomly sampled 100 examples to conduct the user study. The statistical results demonstrate that SiamLayout achieves clear superiority across multiple evaluation dimensions.
E More Details on Datasets and Benchmarks
@@ E.1 LayoutSAM Dataset and Benchmark
E.2 Layout Planning Dataset and Benchmark
@@
F More Analysis on Modal Competition
@@
G More Qualitative Results
We present more qualitative results of SiamLayout-SD3 in Fig. 16. Experimental results show that our proposed method empowers MM-DiT for layout-to-image generation, achieving visually appealing and precisely controllable generation, as demonstrated by the high adherence to complex attributes such as color, texture, and shape.